OpenStudy (anonymous):

Find the slopes of the medians of a triangle whose vertices are P(1, 3), Q(3, 5), and R(6, 2). (Use fraction formats.)

OpenStudy (mertsj):

First you must find the midpoints.

OpenStudy (mertsj):

midpoint of PQ = (2,4), let's call it A slope of median AR = (2-4)/6-2 = -2/4=-1/2